An astronaut on Mars finds that the periodic time of a simple pendulum 1.56 m long is 4.0175 s. Determine the acceleration due to gravity on Mars.


1
Expert's answer
2021-06-29T10:00:20-0400
"T=2\pi\sqrt{\frac{l}{g}}\\4.0175 =2\pi\sqrt{\frac{1.56}{g}}\\g=3.82\frac{m}{s^2}"